# Adaptive Domain Scaling for Personalized Sequential Modeling in Recommenders

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
With the exponential growth of digital contents and the widespread use of the internet, recommender systems have become a vital role for enhancing user experience and alleviating information overload (Zhang et al., 2019). In real-world applications, for improving user retention and promote business benefit, the demands of industrial recommendation are widely distributed cross multiple domains (Jiang et al., 2022; Li et al., 2023). For example, as shown in Fig 1, in Douyin [^1] https://www.douyin.com/, one of the largest video-watching apps in the world, the major domains include Short-videos, Live-preview, and Live-slide, where users can watch short-videos, live-streams, and enjoy the e-commerce and local-life services. Besides, due to its billion-scale user volume, different user groups, like users from different countries, with different genders, highly-active or not, also contribute to different domains. As the data distrbutions are quite diversed across different domains, it poses a significant multi-domain modeling problem for the recommender system (Zhang et al., 2024).
@@
Sequential modeling plays a vital role in industrial recommenders, among which the most popular and effective methods are the target-aware attention based methods, for example, deep interest network (DIN) (Zhou et al., 2018), feature co-action network (CAN) (Bian et al., 2022), and multi-head attention (MHA) (Vaswani et al., 2017). In the advertising system of Douyin, after omitting the target attention module from the conversion-rate prediction (CVR) model, the area under ROC (AUC) metric decreases by over 0.2%, which is a significantly unacceptable loss and substantiates the validity of target attention based sequential modeling. Despite its significance, the impacts of multi-domain discrepancies are less considered in existing target attention methods, leaving a remarkable gap for the area of multi-domain modeling.
Generally speaking, the current target attention mechanism for user sequence can be formulated as a typical query-key-value modeling paradigm: $g(Rep_{cand},Rep_{seq})\times Rep_{seq}$, where $Rep_{cand}$ denotes the representation of the candidate target item whose click/convert probabilities need to be predicted, $Rep_{seq}$ indicates the user sequence embeddings, and $g$ calculates the attention weight between any sequence-item and target-item pairs. As discussed previously, current industrial recommenders generally follow a shared-bottom embedding paradigm, which means that 1) the embedding table of both the candidate item and user behavioral items are fully shared, with no consideration of the distinctions between items and users belonging to different domains, and 2) the candidate item serves as a shared query for different keys/values, with no consideration of the distinctions between the multiple items in user sequence which generally occurs across multiple domains. Accordingly, this poses potential challenges to the current multi-domain recommenders from two aspects:
@@
  • We deploy ADS in both the advertisement system and the e-commercial system of Douyin at ByteDance, which brings significant 1.09% and 0.79% lifts of total revenue in the Douyin ads system and the e-commercial system, respectively. Currently, ADS has been fully deployed in many recommendation systems at ByteDance, serving billions of users.

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
With the exponential growth of digital contents and the widespread use of the internet, recommender systems have become a vital role for enhancing user experience and alleviating information overload (Zhang et al., 2019; Chai et al., 2022; Lu et al., 2025). In real-world applications, for improving user retention and promote business benefit, the demands of industrial recommendation are widely distributed cross multiple domains (Jiang et al., 2022; Li et al., 2023). For example, as shown in Fig 1, in Douyin [^1] https://www.douyin.com/, one of the largest video-watching apps in the world, the major domains include Short-videos, Live-preview, and Live-slide, where users can watch short-videos, live-streams, and enjoy the e-commerce and local-life services. Besides, due to its billion-scale user volume, different user groups, like users from different countries, with different genders, highly-active or not, also contribute to different domains. As the data distrbutions are quite diversed across different domains, it poses a significant multi-domain modeling problem for the recommender system (Zhang et al., 2024).
@@
Sequential modeling plays a vital role in industrial recommenders, among which the most popular and effective methods are the target-aware attention based methods, for example, deep interest network (DIN) (Zhou et al., 2018), feature co-action network (CAN) (Bian et al., 2022), and multi-head attention (MHA) (Vaswani et al., 2017). Despite its significance, the impacts of multi-domain discrepancies are less considered in existing target attention methods, leaving a remarkable gap for the area of multi-domain modeling. Generally speaking, the current target attention mechanism for user sequence can be formulated as a typical query-key-value modeling paradigm: $g(Rep_{cand},Rep_{seq})\times Rep_{seq}$, where $Rep_{cand}$ denotes the representation of the candidate target item whose click/convert probabilities need to be predicted, $Rep_{seq}$ indicates the user sequence embeddings, and $g$ calculates the attention weight between any sequence-item and target-item pairs. As discussed previously, current industrial recommenders generally follow a shared-bottom embedding paradigm, which means that 1) the embedding table of both the candidate item and user behavioral items are fully shared, with no consideration of the distinctions between items and users belonging to different domains, and 2) the candidate item serves as a shared query for different keys/values, with no consideration of the distinctions between the multiple items in user sequence which generally occurs across multiple domains. Accordingly, this poses potential challenges to the current multi-domain recommenders from two aspects:
@@
  • We deploy ADS in both the advertisement system and the e-commercial system of Douyin at ByteDance, which brings significant 1.00% and 0.79% lifts of total revenue in the Douyin ads system and the e-commercial system, respectively. Currently, ADS has been fully deployed in many recommendation systems at ByteDance, serving billions of users.